# RANGER: A Monocular Zero-Shot Semantic Navigation Framework through Contextual Adaptation

### v1

@@ I INTRODUCTION
Efficient visual target localization in complex environments is a core capability of embodied agents, with broad applications in service robotics, human–robot interaction, and autonomous monitoring systems. With the recent development of multimodal foundation models, including large language models (LLMs) (e.g., [OpenAI2023, Liu2023, bai2023qwen, dubey2024llama, team2023gemini]) and open-vocabulary vision models (e.g., [liu2024grounding, kirillov2023segment, cheng2024yolo]), the performance of zero-shot object navigation (e.g., [Kuang2024, Gadre2023, yu2023l3mvn, Wu2024Voronav, yokoyama2024vlfm, yin2025unigoal]) has significantly improved. Such methods allow robots to localize arbitrary objects without task-specific fine-tuning, thereby demonstrating strong potential for deployment in dynamic and previously unseen environments.
However, existing approaches still face major limitations. First, most methods rely heavily on depth sensing and precise pose estimation, usually involving global mapping/reconstruction, which are typically feasible only in simulation or with specialized hardware. In real-world scenarios, sensor noise, calibration errors, data collection, and hardware costs substantially restrict their practicality, similarly to indoor localization scenarios [followus]. Second, current systems have yet to better leverage contextual knowledge available in everyday video streams, which can provide valuable priors for efficient exploration and rapid adaptation to novel environments.
This naturally raises a fundamental question: Can a navigation framework be built relying solely on RGB input? Such a framework would be inherently low-cost, easy to deploy, and capable of directly utilizing videos captured from arbitrary devices for rapid adaptation to new environments. Notably, recent data-driven methods such as MASt3R [leroy2024grounding], Dust3R [wang2024dust3r], Fast3R [yang2025fast3r], and VGGT [wang2025vggt] have demonstrated remarkable performance and generalization in 3D reconstruction tasks, opening new opportunities to achieve efficient navigation with RGB-only input.
@@ II Related Work
Object Goal Navigation. The task of locating a specified object in an unknown environment is widely applicable in robotic systems. Current approaches to object-goal navigation can be categorized into: (1) Approaches involving the use of pre-trained visual encoders to transform first-person observations into visual vectors, which are then processed by a navigation policy network through extensive imitation or supervised learning (e.g.,  [yadav2023ovrl, ramrakhya2023pirlnav, zhang2024navid, batra2020objectnav, ehsani2023spoc, chen2023object, huang2023embodied]), thereby equipping robots with semantic navigation abilities. However, these approaches suffer from overfitting to simulation environments, which limits their generalization to real-world settings. Additionally, they are constrained by predefined object categories, restricting their ability to navigate to arbitrary unseen objects or visual targets. (2) Approaches that focus on Zero-Shot Object Navigation (e.g.,  [Kuang2024, Gadre2023, yokoyama2024vlfm, Wu2024Voronav, yin2025unigoal, yuan2025being0, poliformer2024]). These methods explicitly construct environmental maps and utilize VLMs [Liu2023, OpenAI2023, bai2023qwen] or LLMs [dubey2024llama, team2023gemini] to select the most valuable frontier or waypoint for exploration. For instance, in Cow [Gadre2023], the robot explores the nearest frontier point until the target is detected using CLIP features [clip] and open-vocabulary object detectors [Caron2021]. ESC [zhou2023esc], L3MVN [yu2023l3mvn], and Voronav leverage LLMs for reasoning and decision-making to enhance performance. VLFM [yokoyama2024vlfm] employs a VLM to build semantic values on the map based on first-person observations and textual prompts, selecting the highest-scoring frontier. Their reliance on depth information, precise pose estimation, and task-specific fine-tuning often limits their performance in real-world scenarios, where depth data may be inaccurate or unavailable. These methods also hinder the ability to perform contextual learning, making it challenging for robots to leverage freely available monocular videos to improve exploration efficiency and adapt to new environments.
@@
Monocular 3D Reconstruction. Monocular 3D reconstruction methods have made notable progress recently in enabling the building of 3D maps from single camera inputs. Techniques such as MASt3R [leroy2024grounding], Dust3R [wang2024dust3r], Fast3R [yang2025fast3r], VGGT [wang2025vggt], MASt3R-SLAM [murai2025mast3r-slam], and VGGT-SLAM [vggt-slam] have contributed significantly to this field. However, these methods often rely on passive reconstruction, where the trajectory is pre-collected and does not involve active perception or exploration. This restricts the system’s ability to adapt to dynamic environments and limits its exploration capabilities. Unlike these systems, we aim to incorporate more active, adaptive learning to improve the navigation process without relying on pre-collected trajectories.
@@
In-Context Learning. In-context learning (ICL), first demonstrated in large language models [brown2020language], allows models to perform tasks by leveraging a few examples provided in the input context, without requiring task-specific retraining. ICL has been successfully applied to a variety of tasks, including natural language understanding and object recognition. Recent work has extended this idea to robotics, where robots are trained to perform navigation tasks based on contextual information, such as video or textual prompts. For example, several works have used context from videos or previous maps to train navigation models [zhou2024nolo]. However, these methods often require the target object to appear in the context, which severely limits their practical application, especially in situations where the object is not visible in the initial context. In contrast, our approach does not require the target object to appear, even in the collected offline video. Instead, it enables robots to decide whether to continue exploring or exploit the knowledge contained in the offline video to complete the navigation task. This adaptive approach makes RANGER more versatile, allowing it to handle a broader range of navigation scenarios without the need for predefined object categories or fine-tuning.
@@ IV-A Overview of the Process
RANGER employs a unified architecture that seamlessly handles both zero-shot RGB-only navigation and video-context-augmented navigation without requiring any architectural modifications or retraining. The system maintains a dynamic keyframe bank, which serves as a joint geometric and semantic memory of the environment. Upon receiving each new RGB observation, the backend refines the agent’s pose in real-time and performs loop closure by optimizing relative poses against the stored keyframes. This process incrementally constructs a dense 3D map, enabling metric self-localization. Simultaneously, a semantic module aggregates object features via CLIP [clip] and GroundingDINO [liu2024grounding] into a semantic point cloud, where each point is annotated with category likelihoods and spatial context, creating a queryable semantic memory. To facilitate efficient path planning, a 2D occupancy map is projected from the 3D reconstruction. The hierarchical planner selects high-level waypoints by considering the human-provided goal description, semantic similarity scores of observed regions, and spatial confidence regarding object presence, while a low-level controller translates each waypoint into discrete atomic actions. This perception-planning-action loop operates continuously in real-time until the agent executes the stop action, marking the achievement of high-confidence visual goal localization.
@@ IV-B Online RGB Reconstruction and Localization
To enable metric-aware navigation from monocular RGB inputs, we integrate MASt3R-SLAM [murai2025mast3r-slam], a state-of-the-art real-time dense SLAM system based on the MASt3R two-view 3D reconstruction prior. Its open-source implementation is directly incorporated into RANGER as a plug-and-play geometric backbone.
@@
Global Consistency via Backend Optimization To mitigate drift and ensure long-term consistency, MASt3R-SLAM ensures long-term consistency through second-order global optimization over a pose graph, constructed from sequential and loop-closure edges. Loop candidates are retrieved using ASMK [tolias2013aggregate] on dense visual features and validated via geometric matching.
@@ IV-C Object Point Cloud Merge
We first employ Grounding DINO as an open-vocabulary detector and combine it with Mobile SAM [kirillov2023segment] to extract object masks within the detection boxes of each keyframe stored in the memory bank. Then, using the keyframe confidence masks provided by MASt3R-SLAM, we filter these object regions to obtain high-confidence object point cloud representations in 3D space. For each semantic mask, we further crop the corresponding image region and extract visual features with CLIP. This process ultimately generates a semantic point cloud along with class confidence scores and target category information.
@@ IV-F Path Planning
We use the Fast Marching Method (FMM) [sethian1996fast] as the local planner to generate a path toward the goal. At each step, the robot executes motion commands toward a local subgoal, while the map and target are updated in real time—enabling adaptive, closed-loop navigation in partially observed environments.
@@ V-A Simulation-based Experiment Setup
Navigation in unseen environments. We evaluate our approach on the Habitat HM3Dv2 dataset (HM3D-Semantics-v0.2 [yadav2023habitat], released in the 2023 Habitat Challenge), which includes six goal categories. For our experiments, we select 10 single-floor scenes comprising 279 episodes in total. In each episode, the agent’s starting position is randomly sampled to assess navigation performance in previously unseen environments. Considering the trade-off between system storage and performance, we set the maximum episode length to 300 steps.
@@ V-C Implementation Details
In the Habitat simulation environment, the agent perceives the scene through an RGB camera with a resolution of $640\times 480$ pixels, mounted at a height of 0.88 above the ground plane. For semantic segmentation, we leverage Grounding DINO [liu2024grounding] to generate class-agnostic object proposals and MobileSAM [kirillov2023segment] to produce pixel-accurate masks. Semantic confidence scores for detected objects are computed using CLIP [clip], enabling zero-shot semantic grounding. Our system employs consistent hyper-parameters: keyframe selection threshold $\omega_{k}=0.333$, loop-closure detection threshold $\omega_{l}=0.1$, and robustness regularization weight $\omega_{r}=0.005$. To ensure reliable relocalization, we enforce a stricter geometric consistency check when attaching the current frame to the pose graph: the ratio of inlier feature matches must exceed $0.3$. In addition, the confidence threshold for the point cloud is set to $1.9$.
@@ VI-A Navigation in Unseen Environments
Previously, zero-shot object navigation heavily relied on accurate pose and depth information provided by simulators. To evaluate our approach in more realistic settings, we compare with L3MVN [yu2023l3mvn], a representative work that utilizes LLMs for efficient frontier selection and requires privileged inputs (depth and pose). As detailed in Table I, L3MVN (300) and L3MVN (500) represent performance under step limits of 300 and 500, respectively, both leveraging RGB, depth, and pose data. In contrast, RANGER operates solely on RGB input, without any privileged sensing.
@@ VI-C Ablation Studies
Component Analysis. Table III summarizes the ablation results for two key components: the high-level planner and the object detector. When the high-level planner is removed, performance drops considerably (SR = 34.7%, SPL = 15.3%), and the average steps increase from 172.80 to 202.83. This confirms the critical role of the planner in efficiently guiding the agent toward the target. For the detector comparison, YOLO-World [cheng2024yolo] and DINO yield relatively close results, with DINO showing a slight advantage (SR = 42.7%, SPL = 17.8% vs. 41.7%, 16.7%). This suggests that while replacing the detector has only a modest impact on overall accuracy, stronger visual grounding from DINO provides more consistent navigation performance.

### v2

@@ I INTRODUCTION
Efficient visual target localization in complex environments is a core capability of embodied agents, with broad applications in service robotics, human–robot interaction, and autonomous monitoring systems. With the recent development of multimodal foundation models, including large language models (LLMs) (e.g., [1, 15, 2, 8, 23]) and open-vocabulary vision models (e.g., [16, 12, 7]), the performance of zero-shot object navigation (e.g., [13, 10, 33, 27, 32, 31]) has significantly improved. Such methods allow robots to localize arbitrary objects without task-specific fine-tuning, thereby demonstrating strong potential for deployment in dynamic and previously unseen environments.
However, existing approaches still face major limitations. First, most methods rely heavily on depth sensing and precise pose estimation, usually involving global mapping/reconstruction, which are typically feasible only in simulation or with specialized hardware. In real-world scenarios, sensor noise, calibration errors, data collection, and hardware costs substantially restrict their practicality, similarly to indoor localization scenarios [22]. Second, current systems have yet to better leverage contextual knowledge available in everyday video streams, which can provide valuable priors for efficient exploration and rapid adaptation to novel environments.
This naturally raises a fundamental question: Can a navigation framework be built relying solely on RGB input? Such a framework would be inherently low-cost, easy to deploy, and capable of directly utilizing videos captured from arbitrary devices for rapid adaptation to new environments. Notably, recent data-driven methods such as MASt3R [14], Dust3R [26], Fast3R [30], and VGGT [25] have demonstrated remarkable performance and generalization in 3D reconstruction tasks, opening new opportunities to achieve efficient navigation with RGB-only input.
@@ II Related Work
Object Goal Navigation. The task of locating a specified object in an unknown environment is widely applicable in robotic systems. Current approaches to object-goal navigation can be categorized into: (1) Approaches involving the use of pre-trained visual encoders to transform first-person observations into visual vectors, which are then processed by a navigation policy network through extensive imitation or supervised learning (e.g.,  [28, 20, 37, 3, 9, 6, 11, 34]), thereby equipping robots with semantic navigation abilities. However, these approaches suffer from overfitting to simulation environments, which limits their generalization to real-world settings. Additionally, they are constrained by predefined object categories, restricting their ability to navigate to arbitrary unseen objects or visual targets. (2) Approaches that focus on Zero-Shot Object Navigation (e.g.,  [13, 10, 32, 27, 31, 35, 36]). These methods explicitly construct environmental maps and utilize VLMs [15, 1, 2] or LLMs [8, 23] to select the most valuable frontier or waypoint for exploration. For instance, in Cow [10], the robot explores the nearest frontier point until the target is detected using CLIP features [19] and open-vocabulary object detectors [5]. ESC [39], L3MVN [33], and Voronav leverage LLMs for reasoning and decision-making to enhance performance. VLFM [32] employs a VLM to build semantic values on the map based on first-person observations and textual prompts, selecting the highest-scoring frontier. Their reliance on depth information, precise pose estimation, and task-specific fine-tuning often limits their performance in real-world scenarios, where depth data may be inaccurate or unavailable. These methods also hinder the ability to perform contextual learning, making it challenging for robots to leverage freely available monocular videos to improve exploration efficiency and adapt to new environments.
@@
Monocular 3D Reconstruction. Monocular 3D reconstruction methods have made notable progress recently in enabling the building of 3D maps from single camera inputs. Techniques such as MASt3R [14], Dust3R [26], Fast3R [30], VGGT [25], MASt3R-SLAM [18], and VGGT-SLAM [17] have contributed significantly to this field. However, these methods often rely on passive reconstruction, where the trajectory is pre-collected and does not involve active perception or exploration. This restricts the system’s ability to adapt to dynamic environments and limits its exploration capabilities. Unlike these systems, we aim to incorporate more active, adaptive learning to improve the navigation process without relying on pre-collected trajectories.
@@
In-Context Learning. In-context learning (ICL), first demonstrated in large language models [4], allows models to perform tasks by leveraging a few examples provided in the input context, without requiring task-specific retraining. ICL has been successfully applied to a variety of tasks, including natural language understanding and object recognition. Recent work has extended this idea to robotics, where robots are trained to perform navigation tasks based on contextual information, such as video or textual prompts. For example, several works have used context from videos or previous maps to train navigation models [38]. However, these methods often require the target object to appear in the context, which severely limits their practical application, especially in situations where the object is not visible in the initial context. In contrast, our approach does not require the target object to appear, even in the collected offline video. Instead, it enables robots to decide whether to continue exploring or exploit the knowledge contained in the offline video to complete the navigation task. This adaptive approach makes RANGER more versatile, allowing it to handle a broader range of navigation scenarios without the need for predefined object categories or fine-tuning.
@@ IV-A Overview of the Process
RANGER employs a unified architecture that seamlessly handles both zero-shot RGB-only navigation and video-context-augmented navigation without requiring any architectural modifications or retraining. The system maintains a dynamic keyframe bank, which serves as a joint geometric and semantic memory of the environment. Upon receiving each new RGB observation, the backend refines the agent’s pose in real-time and performs loop closure by optimizing relative poses against the stored keyframes. This process incrementally constructs a dense 3D map, enabling metric self-localization. Simultaneously, a semantic module aggregates object features via CLIP [19] and GroundingDINO [16] into a semantic point cloud, where each point is annotated with category likelihoods and spatial context, creating a queryable semantic memory. To facilitate efficient path planning, a 2D occupancy map is projected from the 3D reconstruction. The hierarchical planner selects high-level waypoints by considering the human-provided goal description, semantic similarity scores of observed regions, and spatial confidence regarding object presence, while a low-level controller translates each waypoint into discrete atomic actions. This perception-planning-action loop operates continuously in real-time until the agent executes the stop action, marking the achievement of high-confidence visual goal localization.
@@ IV-B Online RGB Reconstruction and Localization
To enable metric-aware navigation from monocular RGB inputs, we integrate MASt3R-SLAM [18], a state-of-the-art real-time dense SLAM system based on the MASt3R two-view 3D reconstruction prior. Its open-source implementation is directly incorporated into RANGER as a plug-and-play geometric backbone.
@@
Global Consistency via Backend Optimization To mitigate drift and ensure long-term consistency, MASt3R-SLAM ensures long-term consistency through second-order global optimization over a pose graph, constructed from sequential and loop-closure edges. Loop candidates are retrieved using ASMK [24] on dense visual features and validated via geometric matching.
@@ IV-C Object Point Cloud Merge
We first employ Grounding DINO as an open-vocabulary detector and combine it with Mobile SAM [12] to extract object masks within the detection boxes of each keyframe stored in the memory bank. Then, using the keyframe confidence masks provided by MASt3R-SLAM, we filter these object regions to obtain high-confidence object point cloud representations in 3D space. For each semantic mask, we further crop the corresponding image region and extract visual features with CLIP. This process ultimately generates a semantic point cloud along with class confidence scores and target category information.
@@ IV-F Path Planning
We use the Fast Marching Method (FMM) [21] as the local planner to generate a path toward the goal. At each step, the robot executes motion commands toward a local subgoal, while the map and target are updated in real time—enabling adaptive, closed-loop navigation in partially observed environments.
@@ V-A Simulation-based Experiment Setup
Navigation in unseen environments. We evaluate our approach on the Habitat HM3Dv2 dataset (HM3D-Semantics-v0.2 [29], released in the 2023 Habitat Challenge), which includes six goal categories. For our experiments, we select 10 single-floor scenes comprising 279 episodes in total. In each episode, the agent’s starting position is randomly sampled to assess navigation performance in previously unseen environments. Considering the trade-off between system storage and performance, we set the maximum episode length to 300 steps.
@@ V-C Implementation Details
In the Habitat simulation environment, the agent perceives the scene through an RGB camera with a resolution of $640\times 480$ pixels, mounted at a height of 0.88 above the ground plane. For semantic segmentation, we leverage Grounding DINO [16] to generate class-agnostic object proposals and MobileSAM [12] to produce pixel-accurate masks. Semantic confidence scores for detected objects are computed using CLIP [19], enabling zero-shot semantic grounding. Our system employs consistent hyper-parameters: keyframe selection threshold $\omega_{k}=0.333$, loop-closure detection threshold $\omega_{l}=0.1$, and robustness regularization weight $\omega_{r}=0.005$. To ensure reliable relocalization, we enforce a stricter geometric consistency check when attaching the current frame to the pose graph: the ratio of inlier feature matches must exceed $0.3$. In addition, the confidence threshold for the point cloud is set to $1.9$.
@@ VI-A Navigation in Unseen Environments
Previously, zero-shot object navigation heavily relied on accurate pose and depth information provided by simulators. To evaluate our approach in more realistic settings, we compare with L3MVN [33], a representative work that utilizes LLMs for efficient frontier selection and requires privileged inputs (depth and pose). As detailed in Table I, L3MVN (300) and L3MVN (500) represent performance under step limits of 300 and 500, respectively, both leveraging RGB, depth, and pose data. In contrast, RANGER operates solely on RGB input, without any privileged sensing.
@@ VI-C Ablation Studies
Component Analysis. Table III summarizes the ablation results for two key components: the high-level planner and the object detector. When the high-level planner is removed, performance drops considerably (SR = 34.7%, SPL = 15.3%), and the average steps increase from 172.80 to 202.83. This confirms the critical role of the planner in efficiently guiding the agent toward the target. For the detector comparison, YOLO-World [7] and DINO yield relatively close results, with DINO showing a slight advantage (SR = 42.7%, SPL = 17.8% vs. 41.7%, 16.7%). This suggests that while replacing the detector has only a modest impact on overall accuracy, stronger visual grounding from DINO provides more consistent navigation performance.